# Connecting Jensen–Shannon and Kullback-Leibler Divergences: A New Bound for Representation Learning

### v1

@@ 2 Related work
To overcome these limitations, recent work has introduced variational lower bounds (VLBs) on MI that leverage the expressiveness of deep neural networks. A key example is MINE [3], which leverages the Donsker-Varadhan (DV) representation of the KLD to train a neural network to maximize a lower bound on MI. A related estimator, NWJ [16], arises from the dual representation of $f$-divergences and shares similar objectives. However, theoretical work by McAllester & Stratos, [14] shows that such estimators are fundamentally upper-bounded by $\log(n)$, where $n$ is the number of data samples. Moreover, both MINE and NWJ exhibit high variance, in part due to unstable estimation of the partition function. To tackle this, Poole et al., [20] introduce hybrid VLB methods that generalize existing methods to trade off bias and variance. SMILE [24] further improves upon MINE by clipping the density ratio, which helps reducing the estimator variance. It also introduces a consistency check that exposes weaknesses in existing VLB estimators, especially in high MI regimes, and highlights that optimizing VLBs does not necessarily imply an accurate estimation of a lower bound of MI.
@@
Beyond VLB methods, another class of estimators decouples the estimation and optimization stages. NJEE [22] extends theoretical work by McAllester & Stratos, [14] and estimates MI through entropy subtraction using neural density models. In contrast, DEMI [13] and PCM [26] reformulate MI estimation as a binary classification task: a model is trained to distinguish paired from unpaired samples, and discriminator outputs are then mapped to MI estimates. Interestingly, SMILE [24] can also be interpreted within this discriminator-based framework. As pointed out in [12], its implementation [^1] https://github.com/ermongroup/smile-mi-estimator relies on maximizing an estimator of the JSD [17], with MI subsequently estimated from the outputs of a clipped discriminator. Overall, by disentangling optimization and estimation using a two-step procedure, these methods alleviate some instability issues of VLBs. More recently, Letizia et al., [12] extended this setup by incorporating general $f$-divergences and empirically observed that the JSD (referred as GAN-DIME) provides the best tradeoff between bias and variance, outperforming alternatives based on KLD or Hellinger distance with a relatively low-variance MI estimation. While these two-step estimators are currently among the most accurate MI estimates in practice, they require retraining whenever the underlying joint distribution changes. This limitation is particularly restrictive in representation learning, where the latent variables evolve during optimization. As a result, these methods cannot be reused as plug-in MI estimators during training. In contrast, our bound can be directly optimized via JSD, using a single discriminator, enabling stable, end-to-end training.
@@
Another popular approach is Deep InfoMax from Hjelm et al., [8], which shifts the focus from strictly maximizing MI to maximizing other measures of statistical dependence, particularly the JSD-based objective $\rm I_{\rm JS}$}. Unlike KLD-based bounds used in MINE or NWJ, the JSD is symmetric and bounded, often resulting in more stable optimization, and, unlike CPC and MINE, does not require a large batch size. Deep InfoMax leverages a JSD-based MI estimator introduced by Nowozin et al., [17]. To support the use of JSD-based optimization objective, Hjelm et al., [8] empirically show that for randomly sampled discrete sparse distributions, $\rm I_{\rm JS}$} correlates well with true MI, suggesting it can act as a meaningful proxy. In this work, we go further and provide a theoretical justification: we show that maximizing $\rm I_{\rm JS}$} provably increases a lower bound on the true mutual information.
Jensen-Shannon divergence in GANs: The link between JSD and discriminator training has its roots in Generative Adversarial Networks (GANs), where Goodfellow et al., [6] showed that, under an optimal discriminator, the GAN objective minimizes an affine transformation of the JSD between the real and generated data distributions. This was later generalized in the $f$-GAN framework [17], where GAN training corresponds to a variational lower bound on a general $f$-divergence, estimated via a learned discriminator. Chen et al., [4] proposed InfoGAN, which incorporates MI maximization within the GAN framework using a variational approximation originally introduced by Barber & Agakov, [2]. Rather than estimating the MI between joint and marginal distributions directly, InfoGAN targets the MI between a subset of latent variables and the generator’s output, approximated using an auxiliary network and a variational lower bound from Barber & Agakov, [2]. While InfoGAN brings information-theoretic principles into adversarial training, it does so in a generative setting with a specific structural assumption on the latent space.
@@ 3 Joint range of $f$-divergences
Theorem 3.1 (Harremoës-Vajda, 2011).
@@ 4.1 General lower bound on Kullback–Leibler divergence by Jensen–Shannon divergence
Following Theorem 3.1 by Harremoës & Vajda, [7], this joint range can be fully characterized by restricting $p$, $q$ to be pairs of Bernoulli distributions compatible with JSD and KLD,
@@
In particular, this result holds when $p$ and $q$ are respectively the joint and the product of the marginals, leading to a novel lower bound in information theory between $\rm I_{\rm JS}$} and MI. Theorem 4.1’s proof relies on a conjectured functional property, supported by extensive numerical evidence (see Appendix B.1).
@@ 5.2 Complex Gaussian and non-Gaussian distributions
We follow the protocol of Letizia et al., [12], who benchmark MI estimators in scenarios where the true MI is known analytically. For each estimator, we report the value of the lower bound objective and the ground truth MI. For our method, we additionally report the estimated MI obtained using the procedure described in Section 4.3 and equivalent to GAN-DIME [12].
@@
We also replicate the more challenging benchmark from Letizia et al., [12], following recommendations from Czyż et al., [5] using complex Gaussian transformations (half-cube and asinh, Fig 4) and non-Gaussian distributions (Student and uniform distributions, Fig 4). Across these settings (see details in Appendix E.2), our lower bound and its derived MI estimations remain competitive, with lower variance VLBs, confirming that the proposed JSD-based objective generalizes beyond Gaussian settings.
@@ 5.3 Variational Information Bottleneck Experiments
To assess the practical utility of our JSD-LB in representation learning, we evaluate it within the Information Bottleneck (IB) framework [1, 19], which aims to learn compact yet predictive representations $T$ from inputs $U$ for targets $V$, by maximizing ${\rm I}\left[T;V\right]-{\rm I}\left[T;U\right]$. We reproduce the MNIST experiments of Pan et al., [19], where $U$ denotes an image and $V$ its label, using the same architecture and setup. In our setting (JSD-LB), ${\rm I}\left[T;V\right]$ is optimized by maximizing the JSD via the cross-entropy loss. Following Pan et al., [19], we assess generalization, adversarial robustness, and out-of-distribution (OOD) detection. Robustness is evaluated by computing the average classification accuracy on adversarially perturbed inputs, where pixel-wise pertubation are obtained using a one-step gradient-based attack with perturbation magnitudes $\epsilon\in\{0.1,0.2,0.3\}$. OOD detection is tested on synthetic Gaussian noise samples, using standard OOD detection metrics including True Positive Rate at $95\%$(TPR95), Area Under the ROC Curve (AUROC), Area Under the Precision-Recall Curve (AUPR) and Detection Error.
@@ Acknowledgment
We thank the anonymous reviewers for their insightful comments and helpful suggestions which improved the quality of this paper. R.D. received a Marie Skłodowska-Curie grant No 101154248 (project: SafeREG). This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-SafeREG, 2025-SafeREG). The research leading to these results has received funding from the program "Investissements d’avenir" ANR-10-IAIHU-06. W.W. is supported by the National Institutes of Health (P41EB028741 and R01EB032387).
@@ A.2 Two-step MI estimators
Since MINE suffers from high-variance estimations, Song & Ermon, [24] proposed to clip the outputs of the neural network $T_{\theta}$, which estimates the log-density ratio, when estimating the partition function:
@@
In practice, the implementation of SMILE is significantly different from MINE. Indeed, Song & Ermon, [24] proposed to train the neural network $T_{\theta}$, by maximizing a Jensen-Shannon MI estimator from [17], as in the Deep InfoMax procedure [8]. SMILE is, therefore, a two-step estimator.
@@
$\mathcal{J}_{f}(T)=\mathbb{E}_{(u,v)\sim p_{UV}(u,v)}\left[T(u,v)-f^{*}\left(T\left(x,\sigma\left(y\right)\right)\right)\right]$ (36)
@@
In practice, Letizia et al., [12] proposed to use three $f$ functions that are related to the Kullback–Leibler divergence, GAN, and Hellinger distance squared:
@@ Appendix B Appendix: Proofs of theorems and derivations
Figure 5: Visualization of the negative of the determinant of the Jacobian. We can observe that it is constantly positive, meaning that the determinant of the Jacobian is always negative.
Conjecture B.5.
The determinant of the Jacobian $J$ defined in Eq.(47) is negative on the interior  $\Omega^{\mathrm{o}}=\{(\mu,\nu)\in(0,1)\times(0,1)\mid\nu<\mu\}$ of $\Omega$, i.e.:
$\forall(\mu,\nu)\in\Omega^{\mathrm{o}},\quad\det(J(\mu,\nu))<0\ .$ (52)
While we have not found an analytical proof of the conjecture, it can be verified numerically, as shown in Figure 5. The negative of the determinant of the Jacobian is always positive.
@@
Following the Theorem 3.1 by Harremoës & Vajda, [7], the joint range between the JSD and the KLD can be fully characterized by restricting $p$, $q$ to be pairs of Bernoulli distributions compatible with JSD and KLD,
@@
Using Conjecture B.5, the determinant of the Jacobian is negative for all inner points in $\Omega$. According to the open mapping theorem, the inner points are therefore mapped into interior points of the range. To determine the lower envelope of the image, we examine the image of the boundary $\delta\Omega=(S_{1},S_{2},S_{3})$. Given that, $S_{1}$ and $S_{2}$ respectively map to a single point $(0,0)$ and a horizontal line $y=\infty$, $S_{3}$ defines the lower envelope of the range and is, therefore, our bound.
@@ E.2.2 Analysis for different values of batch size $b$
Our VLB is robust to changes in batch size $b$. In Figures 9, 9, and 9, we observe that MINE and NWJ exhibit high variance, even for large batch sizes (e.g., $b=128$), while JSD-LB and CPC remain more stable. Notably, MINE’s instability can cause substantial overestimation of its lower bound, occasionally exceeding the true MI. CPC, on the other hand, provides low-variance estimates but is upper bounded by the contrastive limit $\log(b)$, which results in higher bias in high-MI regimes.

### v2

@@ 2 Related work
To overcome these limitations, recent work has introduced variational lower bounds (VLBs) on MI that leverage the expressiveness of deep neural networks. A key example is MINE [3], which leverages the Donsker-Varadhan (DV) representation of the KLD to train a neural network to maximize a lower bound on MI. A related estimator, NWJ [16], arises from the dual representation of $f$-divergences and shares similar objectives. However, theoretical work by McAllester and Stratos [14] shows that such estimators are fundamentally upper-bounded by $\log(n)$, where $n$ is the number of data samples. Moreover, both MINE and NWJ exhibit high variance, in part due to unstable estimation of the partition function. To tackle this, Poole et al. [20] introduce hybrid VLB methods that generalize existing methods to trade off bias and variance. SMILE [24] further improves upon MINE by clipping the density ratio, which helps reducing the estimator variance. It also introduces a consistency check that exposes weaknesses in existing VLB estimators, especially in high MI regimes, and highlights that optimizing VLBs does not necessarily imply an accurate estimation of a lower bound of MI.
@@
Beyond VLB methods, another class of estimators decouples the estimation and optimization stages. NJEE [22] extends theoretical work by McAllester and Stratos [14] and estimates MI through entropy subtraction using neural density models. In contrast, DEMI [13] and PCM [26] reformulate MI estimation as a binary classification task: a model is trained to distinguish paired from unpaired samples, and discriminator outputs are then mapped to MI estimates. Interestingly, SMILE [24] can also be interpreted within this discriminator-based framework. As pointed out in [12], its implementation [^1] https://github.com/ermongroup/smile-mi-estimator relies on maximizing an estimator of the JSD [17], with MI subsequently estimated from the outputs of a clipped discriminator. Overall, by disentangling optimization and estimation using a two-step procedure, these methods alleviate some instability issues of VLBs. More recently, Letizia et al. [12] extended this setup by incorporating general $f$-divergences and empirically observed that the JSD (referred as GAN-DIME) provides the best tradeoff between bias and variance, outperforming alternatives based on KLD or Hellinger distance with a relatively low-variance MI estimation. While these two-step estimators are currently among the most accurate MI estimates in practice, they require retraining whenever the underlying joint distribution changes. This limitation is particularly restrictive in representation learning, where the latent variables evolve during optimization. As a result, these methods cannot be reused as plug-in MI estimators during training. In contrast, our bound can be directly optimized via JSD, using a single discriminator, enabling stable, end-to-end training.
@@
Another popular approach is Deep InfoMax from Hjelm et al. [8], which shifts the focus from strictly maximizing MI to maximizing other measures of statistical dependence, particularly the JSD-based objective $\rm I_{\rm JS}$}. Unlike KLD-based bounds used in MINE or NWJ, the JSD is symmetric and bounded, often resulting in more stable optimization, and, unlike CPC and MINE, does not require a large batch size. Deep InfoMax leverages a JSD-based MI estimator introduced by Nowozin et al. [17]. To support the use of JSD-based optimization objective, Hjelm et al. [8] empirically show that for randomly sampled discrete sparse distributions, $\rm I_{\rm JS}$} correlates well with true MI, suggesting it can act as a meaningful proxy. In this work, we go further and provide a theoretical justification: we show that maximizing $\rm I_{\rm JS}$} provably increases a lower bound on the true mutual information.
Jensen-Shannon divergence in GANs: The link between JSD and discriminator training has its roots in Generative Adversarial Networks (GANs), where Goodfellow et al. [6] showed that, under an optimal discriminator, the GAN objective minimizes an affine transformation of the JSD between the real and generated data distributions. This was later generalized in the $f$-GAN framework [17], where GAN training corresponds to a variational lower bound on a general $f$-divergence, estimated via a learned discriminator. Chen et al. [4] proposed InfoGAN, which incorporates MI maximization within the GAN framework using a variational approximation originally introduced by Barber and Agakov [2]. Rather than estimating the MI between joint and marginal distributions directly, InfoGAN targets the MI between a subset of latent variables and the generator’s output, approximated using an auxiliary network and a variational lower bound from Barber and Agakov [2]. While InfoGAN brings information-theoretic principles into adversarial training, it does so in a generative setting with a specific structural assumption on the latent space.
@@ 3 Joint range of $f$-divergences
Theorem 3.1 (Harremoës-Vajda, 2011 [7]).
@@ 4.1 General lower bound on Kullback–Leibler divergence by Jensen–Shannon divergence
Following Theorem 3.1 by Harremoës and Vajda [7], this joint range can be fully characterized by restricting $p$, $q$ to be pairs of Bernoulli distributions compatible with JSD and KLD,
@@
In particular, this result holds when $p$ and $q$ are respectively the joint and the product of the marginals, leading to a novel lower bound in information theory between $\rm I_{\rm JS}$} and MI.
@@ 5.2 Complex Gaussian and non-Gaussian distributions
We follow the protocol of Letizia et al. [12], who benchmark MI estimators in scenarios where the true MI is known analytically. For each estimator, we report the value of the lower bound objective and the ground truth MI. For our method, we additionally report the estimated MI obtained using the procedure described in Section 4.3 and equivalent to GAN-DIME [12].
@@
We also replicate the more challenging benchmark from Letizia et al. [12], following recommendations from Czyż et al. [5] using complex Gaussian transformations (half-cube and asinh, Fig 4) and non-Gaussian distributions (Student and uniform distributions, Fig 4). Across these settings (see details in Appendix E.2), our lower bound and its derived MI estimations remain competitive, with lower variance VLBs, confirming that the proposed JSD-based objective generalizes beyond Gaussian settings.
@@ 5.3 Variational Information Bottleneck Experiments
To assess the practical utility of our JSD-LB in representation learning, we evaluate it within the Information Bottleneck (IB) framework [1, 19], which aims to learn compact yet predictive representations $T$ from inputs $U$ for targets $V$, by maximizing ${\rm I}\left[T;V\right]-{\rm I}\left[T;U\right]$. We reproduce the MNIST experiments of Pan et al. [19], where $U$ denotes an image and $V$ its label, using the same architecture and setup. In our setting (JSD-LB), ${\rm I}\left[T;V\right]$ is optimized by maximizing the JSD via the cross-entropy loss. Following Pan et al. [19], we assess generalization, adversarial robustness, and out-of-distribution (OOD) detection. Robustness is evaluated by computing the average classification accuracy on adversarially perturbed inputs, where pixel-wise pertubation are obtained using a one-step gradient-based attack with perturbation magnitudes $\epsilon\in\{0.1,0.2,0.3\}$. OOD detection is tested on synthetic Gaussian noise samples, using standard OOD detection metrics including True Positive Rate at $95\%$(TPR95), Area Under the ROC Curve (AUROC), Area Under the Precision-Recall Curve (AUPR) and Detection Error.
@@ Acknowledgment
We thank the anonymous reviewers for their insightful comments and helpful suggestions, which improved the quality of this paper. We would also like to thank Yury Polyanskiy for pointing us to the joint range. R.D. received a Marie Skłodowska-Curie grant No 101154248 (project: SafeREG). This work was performed using HPC resources from GENCI–IDRIS (Grant 2024-SafeREG, 2025-SafeREG). The research leading to these results has received funding from the program "Investissements d’avenir" ANR-10-IAIHU-06. W.W. is supported by the National Institutes of Health (P41EB028741 and R01EB032387).
@@ A.2 Two-step MI estimators
Since MINE suffers from high-variance estimations, Song and Ermon [24] proposed to clip the outputs of the neural network $T_{\theta}$, which estimates the log-density ratio, when estimating the partition function:
@@
In practice, the implementation of SMILE is significantly different from MINE. Indeed, Song and Ermon [24] proposed to train the neural network $T_{\theta}$, by maximizing a Jensen-Shannon MI estimator from [17], as in the Deep InfoMax procedure [8]. SMILE is, therefore, a two-step estimator.
@@
$\mathcal{J}_{f}(T)=\mathbb{E}_{(u,v)\sim p_{UV}(u,v)}\left[T(u,v)-f^{*}\left(T\left(u,\sigma\left(v\right)\right)\right)\right]$ (36)
@@
In practice, Letizia et al. [12] proposed to use three $f$ functions that are related to the Kullback–Leibler divergence, GAN, and Hellinger distance squared:
@@ Appendix B Appendix: Proofs of theorems and derivations
Lemma B.5.
The determinant of the Jacobian $J$ defined in Eq. (47) is negative on the interior
$\Omega^{\mathrm{o}}=\{(\mu,\nu)\in(0,1)^{2}:\nu<\mu\},$ (52)
i.e.,
$\forall(\mu,\nu)\in\Omega^{\mathrm{o}},\qquad\det J(\mu,\nu)<0.$ (53)
Proof.
Let $(\mu,\nu)\in\Omega^{\mathrm{o}}$ and define
$x=\frac{\mu}{\nu},\qquad y=\frac{1-\mu}{1-\nu}.$ (54)
Since $0<\nu<\mu<1$, we have
$x>1,\qquad 0<y<1,\qquad x-y=\frac{\mu-\nu}{\nu(1-\nu)}>0.$ (55)
Moreover,
$\frac{1+y}{1+x}=\frac{\nu(2-\mu-\nu)}{(\mu+\nu)(1-\nu)}.$ (56)
Using the formulas established in the lemma B.4, we obtain
$\frac{\partial j}{\partial\mu}(\mu,\nu)$ $=\frac{1}{2}\log\!\left(\frac{\mu(2-\mu-\nu)}{(\mu+\nu)(1-\mu)}\right)=\frac{1}{2}\log\!\left(\frac{x(1+y)}{y(1+x)}\right),$
$\frac{\partial j}{\partial\nu}(\mu,\nu)$ $=\frac{1}{2}\log\!\left(\frac{\nu(2-\mu-\nu)}{(\mu+\nu)(1-\nu)}\right)=\frac{1}{2}\log\!\left(\frac{1+y}{1+x}\right),$
$\frac{\partial k}{\partial\mu}(\mu,\nu)$ $=\log\!\left(\frac{\mu(1-\nu)}{\nu(1-\mu)}\right)=\log\!\left(\frac{x}{y}\right),$
$\frac{\partial k}{\partial\nu}(\mu,\nu)$ $=-\left(\frac{\mu}{\nu}-\frac{1-\mu}{1-\nu}\right)=-(x-y).$
Now set
$A=\log\!\left(\frac{x}{y}\right),\qquad B=\log\!\left(\frac{1+x}{1+y}\right).$ (57)
Since $x>1>y>0$, we have $A>0$ and $B>0$. Also,
$\frac{\partial j}{\partial\mu}=\frac{1}{2}(A-B),\qquad\frac{\partial j}{\partial\nu}=-\frac{1}{2}B,\qquad\frac{\partial k}{\partial\mu}=A,\qquad\frac{\partial k}{\partial\nu}=-(x-y).$ (58)
Hence
$\det J(\mu,\nu)$ $=\frac{1}{2}(A-B)\,[-(x-y)]-\left(-\frac{1}{2}B\right)A$
$=\frac{1}{2}\Bigl(AB-(x-y)(A-B)\Bigr).$ (59)
Therefore, proving $\det J(\mu,\nu)<0$ is equivalent to proving
AB<(xy)(AB),AB<(x-y)(A-B), (60)
or, after dividing by the positive quantity AB(xy)AB(x-y),
$\frac{1}{x-y}<\frac{1}{B}-\frac{1}{A}.$ (61)
Define
$h(t)=\frac{1}{\log\!\left(\frac{x+t}{y+t}\right)},\qquad t\geq 0.$ (62)
Then
$h(1)-h(0)=\frac{1}{B}-\frac{1}{A},$ (63)
so it suffices to show that
$h(1)-h(0)>\frac{1}{x-y}.$ (64)
By the fundamental theorem of calculus,
$h(1)-h(0)=\int_{0}^{1}h^{\prime}(t)\,dt.$ (65)
Thus a sufficient condition for proving $\det J(\mu,\nu)<0$ is to show that
$h^{\prime}(t)>\frac{1}{x-y}\qquad\text{for all }0\leq t\leq 1.$ (66)
A direct computation gives
$h^{\prime}(t)=\frac{x-y}{(x+t)(y+t)\,\log^{2}\!\left(\frac{x+t}{y+t}\right)}.$ (67)
Let
$a=x+t,\qquad b=y+t.$ (68)
Then $a>b>0$ and ab=xya-b=x-y. Condition (66) is equivalent to
$\frac{a-b}{ab\,\log^{2}(a/b)}>\frac{1}{a-b},$ (69)
that is,
$(a-b)^{2}>ab\,\log^{2}\!\left(\frac{a}{b}\right).$ (70)
Writing $z=a/b>1$, this becomes
$(z-1)^{2}>z(\log z)^{2},$ (71)
or equivalently
$\frac{z-1}{\sqrt{z}}>\log z.$ (72)
To prove (72), let $r=\frac{1}{2}\log z>0$, so that $z=e^{2r}$. Then
$\frac{z-1}{\sqrt{z}}=\frac{e^{2r}-1}{e^{r}}=e^{r}-e^{-r}=2\sinh r,\qquad\log z=2r.$ (73)
Hence (72) is equivalent to
$\sinh r>r,$ (74)
which is true for every $u>0$ since
$\sinh r-r=\sum_{n\geq 1}\frac{r^{2n+1}}{(2n+1)!}>0.$ (75)
Therefore $h^{\prime}(t)>\frac{1}{x-y}$ for all $t\geq 0$, and so
$h(1)-h(0)=\int_{0}^{1}h^{\prime}(t)\,dt>\int_{0}^{1}\frac{dt}{x-y}=\frac{1}{x-y}.$ (76)
This proves (61), hence (60), and finally, by (59),
$\det J(\mu,\nu)<0.$ (77)
Therefore, for every $(\mu,\nu)\in\Omega^{\mathrm{o}}$, the Jacobian determinant is strictly negative. ∎
@@
Following the Theorem 3.1 by Harremoës and Vajda [7], the joint range between the JSD and the KLD can be fully characterized by restricting $p$, $q$ to be pairs of Bernoulli distributions compatible with JSD and KLD,
@@
Using Lemma B.5, the determinant of the Jacobian is negative for all inner points in $\Omega$. According to the open mapping theorem, the inner points are therefore mapped into interior points of the range. To determine the lower envelope of the image, we examine the image of the boundary $\delta\Omega=(S_{1},S_{2},S_{3})$. Given that, $S_{1}$ and $S_{2}$ respectively map to a single point $(0,0)$ and a horizontal line $y=\infty$, $S_{3}$ defines the lower envelope of the range and is, therefore, our bound.
@@ E.2.2 Analysis for different values of batch size $b$
Our VLB is robust to changes in batch size $b$. In Figures 8, 8, and 8, we observe that MINE and NWJ exhibit high variance, even for large batch sizes (e.g., $b=128$), while JSD-LB and CPC remain more stable. Notably, MINE’s instability can cause substantial overestimation of its lower bound, occasionally exceeding the true MI. CPC, on the other hand, provides low-variance estimates but is upper bounded by the contrastive limit $\log(b)$, which results in higher bias in high-MI regimes.